# Direct Imitation Learning-based Visual Servoing using the Large Projection Formulation

### v1

@@ 7 Conclusion
In this paper, we have presented Direct Imitation Learning Visual Servoing (D-ILVS), a dynamical system-based imitation learning approach for direct visual servoing. The proposed framework overcomes several limitations of existing approaches. D-ILVS exploits off-the-shelf deep learning-based perception to extract features from raw camera images, augmented with imitation learning layers that generate complex robot trajectories. A key difference from end-to-end learning approaches is that D-ILVS exploits a control theoretical framework to ensure convergence to a given target. The approach has been extensively evaluated with real robot experiments, and compared with two baselines showing superior performance.
Figure 13: Two final image frames captured by the robot camera showing that D-ILVS successfully drives the robot to drop the object into the cup in different cluttered environments.

### v2

@@ 7 Conclusion
In this paper, we have presented Imitation Learning-based Direct Visual Servoing (IL-DVS), a dynamical system-based imitation learning approach for direct visual servoing. The proposed framework overcomes several limitations of existing approaches. IL-DVS exploits off-the-shelf deep learning-based perception to extract features from raw camera images, augmented with imitation learning layers that generate complex robot trajectories. A key difference from end-to-end learning approaches is that IL-DVS exploits a control theoretical framework to ensure convergence to a given target. The approach has been extensively evaluated with real robot experiments, and compared with two baselines showing superior performance.